# Moments and the Basic Equilibrium Equations

### v3

@@ 2. Moments of a Weighted System
We begin with a brief overview of affine geometry; see [8] for a thorough exposition. Let $(\mathbb{A},\overrightarrow{\mathbb{A}},\tau)$ be an Euclidean affine space, that is, a triple formed by a point set $\mathbb{A}$, a vector space $\overrightarrow{\mathbb{A}}$ over a field $K$ endowed with an inner product $\cdot$ , and a free action $\tau$ of the additive group of $\overrightarrow{\mathbb{A}}$ on $\mathbb{A}$. The vector $\overrightarrow{v}\in\overrightarrow{\mathbb{A}}$ acts on the point $x\in\mathbb{A}$ by translating it to a point $y\in\mathbb{A}$ and is unique in that sense, so we write
@@
The pair $(\overrightarrow{\mathbb{A}},\tau)$ is the affine structure of the affine space $(\mathbb{A},\overrightarrow{\mathbb{A}},\tau)$, which is often denoted simply by $\mathbb{A}$. A basic and very useful result is the Chasles relation: $\overrightarrow{xy}+\overrightarrow{yz}=\overrightarrow{xz}$.
Let $X$ be a finite subset of an Euclidean affine space $\mathbb{A}$ and $w$ be a function from $X$ to $K$, the field of scalars of $\mathbb{A}$. We refer to $X$ as a configuration of points in $\mathbb{A}$, and to $w$ as a weight function on $X$. Three key maps can be assigned to the weighted system formed by $X$ and $w$, i.e., to the pair $(X,w)$.
@@
The total weight $\mu_{0}$, the first moment $\mu_{1}:\mathbb{A}\longrightarrow\overrightarrow{\mathbb{A}}$ and the second moment $\mu_{2}:\mathbb{A}\longrightarrow K$ of the weighted system $(X,w)$ are respectively given by
$\mu_{0}=\sum_{x\in X}w(x),\ \ \ \mu_{1}(p)=\sum_{x\in X}w(x)\overrightarrow{px},\ \ \text{and}\ \ \mu_{2}(p)=\sum_{x\in X}w(x)\overrightarrow{px}^{2},$
@@
The total weight can be thought of as the “zeroth moment”, hence its notation. It is sometimes convenient to see $\mu_{0},\mu_{1}$ and $\mu_{2}$ as (linear) maps on the weight function.
We state some basic results regarding the moments $\mu_{0}$, $\mu_{1}$ and $\mu_{2}$ of a fixed weighted system $(X,w)$ in $\mathbb{A}$.
@@
Given $p,q\in\mathbb{A}$, we have that $\mu_{1}(p)-\mu_{1}(q)=\mu_{0}\overrightarrow{pq}.$
@@
where $q$ is an arbitrary point of $\mathbb{A}$. We call $G$ the barycenter of $(X,w)$ and denote it by $\text{bar}(X,w)$. Using the barycenter makes it possible to simplify the expressions of the first and second moments, as shown by proposition 2.3 and the Huygens-Leibniz-König theorem below.
@@
Part (1) follows directly from lemma 2.2 and the definition of $G$. For part (2), note that $\mu_{1}$ is constant according to lemma 2.2, so we have that $\mu_{1}(p)=\mu_{1}(x)$. But the total weight of the weighted system $\left(X\setminus\{x\},w\big|_{X\setminus\{x\}}\right)$ is $w(x)\neq 0$ and its first moment at the point $x$ is $\mu_{1}(x)$. So, by applying part (1) to the latter system with $p=x$, we obtain part (2). ∎
Part (2) of proposition 2.3 has a close relation to barycentric coordinates, see remark 4.4. We have the following immediate consequence of lemma 2.2 and proposition 2.3.
@@
For every $p,q\in\mathbb{A}$, we have that
@@
    if $\mu_{0}\neq 0$ and $G=\text{\emph{bar}}(X,w)$, then $\mu_{2}(p)-\mu_{2}(G)=\mu_{0}\,\overrightarrow{pG}^{2}$, for all $p\in\mathbb{A}$,
@@
    if $\mu_{0}=0$ and $\mu_{1}=\overrightarrow{v}$, then $\mu_{2}(p)-\mu_{2}(q)=2\overrightarrow{v}\cdot\overrightarrow{qp},$ for all $p,q\in\mathbb{A}$, and
@@
by $w(x)$, add for all $x\in X$, and apply lemma 2.2. Picking $q=G$ gives (1), and taking $q\in\mathbb{A}$ arbitrary produces (2). To prove (3), set $p=y\in X$. From (2.1), we can write
@@
By lemma 2.2, we can replace $\mu_{1}(y)$ by $\mu_{0}\overrightarrow{yq}+\mu_{1}(q)$ in the latter summation. Leibniz’s identity follows.
The HLK theorem has a broad range of applications. Below we state just two of its corollaries, the second of which is consequential to a significant part of this work, including the study of configurations described in terms of mutual distances, see appendix A.
@@
$\sum_{\{x,y\}\subset X}w(x)w(y)\overrightarrow{xy}^{2}=0$
In appendix B we discuss an extension of corollary 2.5.
@@
Corollary 2.6 has the significant implications below.
@@
We have that $\mu_{1}=\overrightarrow{O}$ if and only if $\mu_{0}=0$ and either of the equivalent equations
(2.2) $\sum_{z\in X}w(z)(\overrightarrow{xz}^{2}-\overrightarrow{yz}^{2})=0,\quad\text{or}\quad\sum_{z\in X\setminus\{x\}}w(z)(\overrightarrow{xy}^{2}-\overrightarrow{yz}^{2}+\overrightarrow{zx}^{2})=0$
@@ 3. Equilibrium Equations
Let $X$ be a finite set of points representing the positions of interacting particles. Let us assume that the force of interaction between two distinct particles $x,y\in X$ can be expressed as
(3.1) $\quad\overrightarrow{F}_{x,y}=\varphi_{x,y}\overrightarrow{xy},\quad\text{with }\ \varphi_{x,y}:\prod_{z\in X}\mathbb{A}\longrightarrow K\ \text{arbitrary}.$
Notice that the expression (3.1) is commonly adopted to describe internal forces in Classical Mechanics, together with conditions on the $\varphi_{x,y}$ in order to comply the homogeneity and isotropy of space [6]. For our purposes, though, it is necessary neither to make these additional assumptions, nor to impose the validity of Newton’s third law of motion, which corresponds to equalities $\varphi_{x,y}=\varphi_{y,x}$ for all $x,y\in X$.
The total force on particle $x$ is $\overrightarrow{F}_{x}=\sum_{y\neq x}\overrightarrow{F}_{x,y}$. As usual, we define an equilibrium of a system of interacting particles as a configuration for which the total force on each particle is zero. For each $x\in X$, let $(X,w_{x})$ be the weighted system defined by
@@
It is clear that each system $(X,w_{x})$ has zero total weight and (constant) first moment $\mu_{1}=\overrightarrow{F}_{x}$. The next result is a direct consequence of the present discussion.
@@
Theorem 3.1 represents a landmark in our work as it bridges the theory of moments of weighted systems of section 2 with applications to equilibrium problems. For instance, theorems 3.1, corollary 2.7 and the extended Leibniz identities of appendix B allow us to write equations for equilibria which are homogeneous on the force coefficients $\varphi_{x,y}$ and have a relatively simple form.
@@
A configuration $X$ is an equilibrium of a systems of particles which interact according to equations (3.1) if and only if, for every pair of points $x,y\in X$, there holds one of the following identities
@@
(3.4) $\sum_{\{z,w\}\subset X}\varphi_{x,z}\varphi_{y,w}\overrightarrow{zw}^{2}=0.$
@@
In section 6 we find another set of equilibrium equations for systems which respect Newton’s third law.
The following remarks contain comments on possible applications.
@@
If a function $f$ depends solely on the mutual distances between particles, then the gradient of $f$ for each individual particle has the form (3.1), so corollary 3.2 can be applied to study the critical points of $f$. In Physics, a force field is called conservative when it is the gradient of a potential function. If each $\varphi_{x,y}$ depends only on the distance from $x$ to $y$, that is, if the force fields $F_{x}:\overrightarrow{F}_{x,y}$, $x,y\in X$ are central according to the definition in [6], then $F_{x}$ is shown to be conservative. If, in addition, all the forces derive from a single potential function, a simple calculation shows that Newton’s third law of motion is valid.
@@
External forces (due to a moving reference frame, for instance) acting on the particles may be dealt with by the above method as long as the total force on each particle remains the summation of terms of the form (3.1). The example of central configurations of nonzero masses and nonzero total mass illustrates this possibility, see section 5, equations (5.3) and (5.4) (notice that there, in view of Newton’s second law of motion, we found it more convenient to work with accelerations instead of forces). Reference [7] discusses related problems.
4. The space of weighted systems with $\mu_{1}={\overrightarrow{O}}$
The previous section motivates a deeper study of weighted systems with zero total weight and zero first moment or, more briefly, weighted systems with $\mu_{1}=\overrightarrow{O}$, cf. corollary 2.4. We leave applications and the relation with the second moment respectively to the next section and the appendices.

### v4

@@ 2. Moments of a Weighted System
We begin with an overview of affine geometry; see [6] for a thorough exposition. Let $(\mathbb{A},\overrightarrow{\mathbb{A}},\tau)$ be an affine space over a field $K$, that is, a triple formed by a point set $\mathbb{A}$, a vector space $\overrightarrow{\mathbb{A}}$ over $K$ and a free action $\tau$ of the additive group of $\overrightarrow{\mathbb{A}}$ on $\mathbb{A}$. The vector $\overrightarrow{v}\in\overrightarrow{\mathbb{A}}$ acts on the point $x\in\mathbb{A}$ by translating it to a point $y\in\mathbb{A}$ and is unique in that sense, so we write
@@
The pair $(\overrightarrow{\mathbb{A}},\tau)$ is the affine structure of the affine space $(\mathbb{A},\overrightarrow{\mathbb{A}},\tau)$, which is often denoted simply by $\mathbb{A}$. If $\overrightarrow{\mathbb{A}}$ is endowed with an inner product, we call $\mathbb{A}$ an Euclidean affine space and use a $\cdot$ to indicate the inner product. A basic and very useful result is the Chasles relation: $\overrightarrow{xy}+\overrightarrow{yz}=\overrightarrow{xz}$.
Let $X$ be a finite subset of an affine space $\mathbb{A}$ and $w$ be a function from $X$ to $K$, the field of scalars of $\mathbb{A}$. We refer to $X$ as a configuration of points in $\mathbb{A}$, and to $w$ as a weight function on $X$. Three key maps can be assigned to the weighted system formed by $X$ and $w$, i.e., to the pair $(X,w)$.
@@
The total weight $\mu_{0}$, the first moment $\mu_{1}:\mathbb{A}\longrightarrow\overrightarrow{\mathbb{A}}$, and (for Euclidean affine spaces) the second moment $\mu_{2}:\mathbb{A}\longrightarrow K$ of the weighted system $(X,w)$ are respectively given by
$\mu_{0}=\sum_{x\in X}w(x),\ \ \ \mu_{1}(p)=\sum_{x\in X}w(x)\overrightarrow{px},\ \ \text{and}\ \ \mu_{2}(p)=\sum_{x\in X}w(x)\overrightarrow{px}^{2},\ \ \forall p\in\mathbb{A},$
@@
The total weight can be thought of as the “zeroth moment”, hence its notation. It is sometimes convenient to see $\mu_{0},$ $\mu_{1}$ and $\mu_{2}$ as linear maps on the weight function (see section 4).
We state some basic results regarding the moments of a fixed weighted system $(X,w)$ in a given affine space $\mathbb{A}$.
@@
For all $p,q\in\mathbb{A}$, we have that $\mu_{1}(p)-\mu_{1}(q)=\mu_{0}\overrightarrow{pq}.$
@@
where $q$ is an arbitrary point of $\mathbb{A}$. We call $G$ the barycenter of $(X,w)$ and denote it by $\text{bar}(X,w)$.
@@
Part (1) follows directly from lemma 2.2 and the definition of barycenter. For part (2), note that $\mu_{1}$ is constant according to lemma 2.2, so we have that $\mu_{1}(p)=\mu_{1}(x)$. But the total weight of the weighted system $\left(X\setminus\{x\},w\big|_{X\setminus\{x\}}\right)$ is $w(x)\neq 0$, and the first moment of $\left(X\setminus\{x\},w\big|_{X\setminus\{x\}}\right)$ at the point $x$ is $\mu_{1}(x)$. So, by applying part (1) to the latter system, with $p=x$, we obtain part (2). ∎
The following are immediate consequences of lemma 2.2 and proposition 2.3.
@@
Corollary 2.5.
If $\mu_{1}=\overrightarrow{O}$, then for every $x\in X$ with $w(x)\neq 0$, we have that $x=\text{bar}\left(X\setminus\{x\},w\big|_{X\setminus\{x\}}\right)$.
Notice that, if $X\setminus\{x\}$ in corollary 2.5 is a simplex with the same dimension as $X$, then $w\big|_{X\setminus\{x\}}$ defines homogeneous barycentric coordinates for $x$ with respect to $X\setminus\{x\}$. This remark is key to some relevant applications and will be explored in detail in a subsequent work.
@@
Let $\mathbb{A}$ be an Euclidean affine space, $(X,w)$ a weighted system in $\mathbb{A}$, and $\mu_{0},\mu_{1},\mu_{2}$ the moments of $(X,w)$. For every $p,q\in\mathbb{A}$, we have that
@@
    if $\mu_{0}\neq 0$ and $G=\text{\emph{bar}}(X,w)$, then $\mu_{2}(p)-\mu_{2}(G)=\mu_{0}\,\overrightarrow{pG}^{2}$, for all $p\in\mathbb{A}$;
@@
    if $\mu_{0}=0$ and $\mu_{1}=\overrightarrow{v}$, then $\mu_{2}(p)-\mu_{2}(q)=2\overrightarrow{v}\cdot\overrightarrow{qp},$ for all $p,q\in\mathbb{A}$; and
@@
by $w(x)$, add for all $x\in X$, and apply lemma 2.2. Picking $q=G$ gives (1), and taking $q\in\mathbb{A}$ arbitrary produces (2). To prove (3), set $p=y\in X$. From equation (2.1), we can write
@@
By lemma 2.2, we can replace $\mu_{1}(y)$ by $\mu_{0}\overrightarrow{yq}+\mu_{1}(q)$ in the latter summation. Leibniz’s identity follows. ∎
The HLK theorem has a broad range of applications. Below we state two of its corollaries, the second of which is consequential to a significant part of this work, including the formulation of equilibrium problems (section 3) and the study of configurations described in terms of mutual distances (section 5 and appendix A).
@@
$\sum_{\{x,y\}\subset X}w(x)w(y)\overrightarrow{xy}^{2}=0.$
@@
Corollary 2.7 has the useful implication below.
@@
The weighted system $(X,w)$ has $\mu_{1}=\overrightarrow{O}$ if and only if $\mu_{0}=0$ and
(2.2) $\sum_{z\in X}w(z)(\overrightarrow{xz}^{2}-\overrightarrow{yz}^{2})=0\quad\text{or}\quad\sum_{z\in X\setminus\{x\}}w(z)(\overrightarrow{xy}^{2}-\overrightarrow{yz}^{2}+\overrightarrow{zx}^{2})=0,$
@@
The relationship between corollaries 2.6 and 2.7 will be clarified in proposition 4.3.
@@ 3. Equilibrium Equations
Let $X\subset\mathbb{A}$ represent a finite set of interacting particles. Let us assume that the force of interaction between two distinct particles $x,y\in X$ can be expressed as
(3.1) $\quad\overrightarrow{F}_{x,y}=\varphi_{x,y}\overrightarrow{xy}.$
The coefficients $\varphi_{x,y}$ are (rather arbitrary) elements of $K$ and may, by allowing $X$ to vary, be functions of the positions of the points in $X$.
The total force on particle $x$ is $\overrightarrow{F}_{x}=\sum_{y\neq x}\overrightarrow{F}_{x,y}$. As customary, we define an equilibrium of a system of interacting particles as a configuration for which the total force on each particle is zero. For each $x\in X$, let $(X,w_{x})$ be the weighted system defined by
@@
It is clear that each system $(X,w_{x})$ has zero total weight and (constant) first moment $\mu_{1}=\overrightarrow{F}_{x}$. The next result is a direct consequence of this key observation.
@@
Theorem 3.1 represents a landmark in our work as it bridges the theory of moments from section 2 with applications to equilibrium problems. For instance, corollary 2.8, theorem 3.1 and proposition 4.3 allow us to write (homogeneous) equations for equilibria which depend only on the force coefficients $\varphi_{x,y}$ and squared mutual distances, and have a relatively simple form.
@@
A configuration $X$ is an equilibrium of a system of particles in an Euclidean affine space which interact according to equations (3.1) if and only if there holds one of the following sets of identities
@@
(3.4) $\sum_{\{u,z\}\subset X}\varphi_{x,u}\varphi_{y,z}\overrightarrow{uz}^{2}=0.$
for every pair of points $x,y\in X$.
@@
Due to their relevance and relation to classical results, we refer to equations (3.3) and (3.4) as the generalized Albouy-Chenciner equations and extended Leibniz identities, respectively. A somewhat curious fact regarding corollary 3.2 is that the Euclidean structure of the affine space need not coincide with the one used to define the distances that may appear in the arguments of $\varphi_{x,y}$.
The following remarks contain comments on applications.
@@
Expression (3.1) is commonly adopted to describe internal forces in Classical Mechanics, together with conditions on the $\varphi_{x,y}$ in order to comply with the homogeneity and isotropy of space [4]. For our purposes, though, it is necessary neither to make these additional assumptions, nor to require the validity of Newton’s third law of motion, which corresponds to equalities $\varphi_{x,y}=\varphi_{y,x}$ for all $x,y\in X$.
@@
If a function $f$ depends solely on the mutual distances between particles, then the gradient of $f$ for each individual particle has the form (3.1), so theorem 3.1 and corollary 3.2 can be applied to study the critical points of $f$. In Physics, a force field is called conservative when it is the gradient of a (potential) function. If each $\varphi_{x,y}$ depends only on the distance from $x$ to $y$, that is, if the force fields $F_{x}:\overrightarrow{F}_{x,y}$, $x,y\in X$ are central according to the definition in [4], then $F_{x}$ is shown to be conservative. In addition, if all the forces derive from a single potential function, a simple calculation shows that Newton’s third law of motion is valid for the corresponding system.
Remark 3.5.
@@
External forces acting on the particles (due to a moving reference frame, for instance) may be dealt with by the above method as long as the total force on each particle remains the summation of terms of the form (3.1). The example of central configurations of nonzero masses and nonzero total mass illustrates this possibility, see section 6, equations (6.3) and (6.4) (notice that there, in view of Newton’s second law of motion, we found it more convenient to work with accelerations instead of forces). Reference [5] discusses related problems.
4. Matrix Formulations
In the next paragraphs we employ basic Linear Algebra to deepen our analysis of the relationship between the moments $\mu_{0},$ $\mu_{1}$ and $\mu_{2}$. The results of this section yield the equilibrium equations (3.4) in section 3, and will be used in section 5 to study the geometry of configurations.
For convenience, from now on we let $X=\{x_{1},\ldots,x_{n}\}$. A weight function $w:X\longrightarrow K$ shall be represented by the weight vector $W\in K^{n}$ whose coordinates are $w(x_{j})=w_{j}$, $j=1,\ldots,n$.
We select a reference frame and introduce affine coordinates in the affine closure of $X$, $\text{Aff}(X)$, in the usual way. Let $\mathbb{W}(X)=K^{n}$ denote the space of weight vectors defined on $X$. We call $\mathbb{W}(X)$ the weight space of $X$. Let us define the $1\times n$ matrix ${\bm{1}}=\left[1\ldots 1\right]$, and the matrices
$\mathcal{X}=\left[x_{ij}\right]_{1\leq j\leq n}\quad\text{and}\quad\mathcal{B}(X)=\left[s_{ij}\right]_{1\leq i,j\leq n},$
where $x_{ij}$ is the $i^{\text{th}}$ coordinate of $x_{j}$ and $s_{ij}=\overrightarrow{x_{i}x_{j}}^{2}$. The matrices $\bm{1}$}, $\mathcal{X}$ and $\mathcal{B}(X)$ establish linear maps on $\mathbb{W}(X)$ corresponding, respectively, to the moments $\mu_{0},\mu_{1}$ and $\mu_{2}$.
Notice that the condition $\mu_{0}=0$ corresponds to the weight vector $W$ belonging to the kernel of $\bm{1}$. Denote $\text{Ker}({\bm{1}})$ by $\mathbb{D}$. The condition $\mu_{1}=\overrightarrow{O}$ corresponds to $\bm{1}W=0$ and $\mathcal{X}W=\bm{0}^{T}$, where $\bm{0}$ is the $1\times n$ matrix of zeros. So $W$ is a weight vector in the kernel of the augmented configuration matrix
$\mathcal{X}_{a}=\left[\begin{array}[]{c}\bm{1}\\ \mathcal{X}\end{array}\right].$
Henceforth we write $\text{Ker}(\mathcal{X}_{a})=\mathbb{W}_{0}(X)$. It is clear that $\mathbb{W}_{0}(X)\leq\mathbb{D}$. Now consider $\mathcal{B}(X)$ as a map on $\mathbb{D}$. According to corollary 2.7, equivalence between sentences (i) and (iii), we have that
$W\in\mathbb{W}_{0}(X)\quad\Longleftrightarrow\quad\exists w_{0}\in K\ \ \text{such that}\ \ \mathcal{B}(X)W=w_{0}\bm{1}^{T}.$
So $w_{0}{\bm{1}}^{T}+\mathcal{B}(X)W={\bm{0}}^{T}$, and we may write
(4.1) $w\in\mathbb{W}_{0}(X)\quad\Longleftrightarrow\quad\mathcal{C}(X)\widetilde{W}={\bm{0}},$
where $\widetilde{W}=(-w_{0},w_{1},\ldots,w_{n})$ and
$\mathcal{C}(X)=\left[\begin{array}[]{cc}0&{\bm{1}}\\ {\bm{1}}^{T}&\mathcal{B}(X)\end{array}\right],$
The determinant of $\mathcal{C}(X)$ is known as the Cayley-Menger determinant of $X$ [6, 7].
Theorem 4.1.
For any configuration $X$, we have that $\emph{Ker}\,\,\mathcal{B}(X)\cap\mathbb{D}\leq\mathbb{W}_{0}(X)\cong\emph{Ker}\,\,\mathcal{C}(X)$, with $\emph{Ker}\,\,\mathcal{B}(X)\cap\mathbb{D}=\mathbb{W}_{0}(X)$ if and only if for every weighted system $(X,w)$ with $\mu_{0}=0$ and $\mu_{1}=\overrightarrow{O}$, we also have $\mu_{2}=0$.
Proof.
If $W\in\text{Ker}\,\,\mathcal{B}(x)\cap\mathbb{D}$, then the restriction of $\mu_{2}$ to $X$ vanishes identically. Thus $\mu_{1}=\overrightarrow{O}$, by corollary 2.7, so $\text{Ker}\,\,\mathcal{B}(x)\cap\mathbb{D}\leq\mathbb{W}_{0}(X)$. The isomorphism $\mathbb{W}_{0}(X)\cong\text{Ker}\,\,\mathcal{C}(X)$ is clear from (4.1). ∎
We notice that the aforementioned natural isomorphism $W\mapsto\widetilde{W}$ in theorem 4.1 seems to have firstly appeared in corollary 3.4 of reference [11].
A configuration $X$ is called co-spherical configuration if there exists a point $p\in\mathbb{A}$ and a scalar $R\in K$ such that $\overrightarrow{px_{i}}^{2}=R^{2}$, for all $i=1,\ldots,n$.
Corollary 4.2.
If $X$ is a co-spherical configuration, then $\emph{Ker}\,\,\mathcal{B}(X)\cap\mathbb{D}=\mathbb{W}_{0}(X)$.
Proof.
Let $p\in\mathbb{A}$, $R\in K$ be as in the definition of co-spherical configuration, and let $W\in\mathbb{W}_{0}(X)$. Since $W\in\mathbb{D}$, we have that
$\mu_{2}(p)=\sum_{i=1}^{n}w_{i}\overrightarrow{px_{i}}^{2}=\mu_{0}R^{2}=0,$
Thus $\mu_{2}$ is the null function, and $W\in\text{Ker}\,\,\mathcal{B}(X),$ by theorem 4.1. ∎
The matrix $\mathcal{B}(X)$ is known as the relative configuration matrix. It is noteworthy that $\mathcal{B}(X)$ plays a key role in the reduction of the $n$-body problem by rotations and translations [2, 21], where it is viewed as a symmetric bilinear form on $\mathbb{D}$. We verify the basic properties of this bilinear form below.
Proposition 4.3.
Let $X$ be a configuration in an affine Euclidean space, and let $\mathcal{B}(X)$ be its relative configuration matrix viewed as a bilinear form on the space $\mathbb{D}$ of weight vectors with $\mu_{0}=0$.
  • (i)
    The quadratic form defined by $\mathcal{B}(X)$ is negative semi-definite.
  • (ii)
    The kernel of $\mathcal{B}(X)$ is $\mathbb{W}_{0}(X)$. In particular, the restriction of $\mathcal{B}(X)$ to $\mathbb{W}_{0}(X)$ is the null form.
Proof.
Item (i) is a direct consequence of the Leibniz identity, see the HLK theorem, part (3).
Notice item (ii) follows directly from corollary 2.6. We provide an alternative proof that helps to clarify the relationship between corollaries 2.6 and 2.7. Consider the basis of $\mathbb{D}$ consisting of the weight vectors $E_{12}=(1,-1,0,\ldots,0),\ldots,E_{1n}=(1,0,\ldots,0,-1)$. By corollary 2.7, $W\in\mathbb{W}_{0}(X)$ if and only if $\mu_{2}(x_{1})=\mu_{2}(x_{i})$, $i=1,\ldots,n,$ that is, if and only if
$E_{1i}^{T}\mathcal{B}(X)W=0,\quad\text{for all}\quad i=2,\ldots,n,$
which is equivalent to $W^{T}\mathcal{B}(X)={\bm{0}}$ on $\mathbb{D}$. ∎
Remark 4.4.
Early on in the next section we shall define the dimension of a configuration. Proposition 4.3(ii) and lemma 5.1 imply that the rank of $\mathcal{B}(X)$ coincides with the dimension of $X$. This fact and proposition 4.3(i) are consistent with $X$ being irreducibly isometrically embedded in $\mathbb{A}$ (see the corollary on page 107 of [7]).
5. Constraints for Mutual Distances
In this section we draw some consequences of theorem 4.1 for describing configurations in terms of mutual distances.